The mass of a dust particle is 7.53 x 10^-10 kg. If there were 6 x 10^4 particles in a giant dust ball on your floor, what would be the mass of that dust ball? Express your answer in scientific notation.
Answer by checkley77(12844)

(7.53 x 10^-10)(6 x 10^4 )=45.18*10^(4-10)=45.18*10^(-6)